OpenStudy (anonymous):

find the slope of line containing pairs of points (-6,-10) (-12,-11)

OpenStudy (anonymous):

(y2-y1)/(x2-x1)

hartnn (hartnn):

the slope of the line containing points(x1,y1)and(x2,y2) is m=(y2-y1)/(x2-x1)

hero (hero):

Insert each point into the formula y = mx + b to get -10 =-6m + b -11 = -12m + b Isolate b in each equation and solve for m

OpenStudy (anonymous):

so -10= -6m -11= -12m

OpenStudy (anonymous):

-4.1

hero (hero):

Isolate b in both equations to get: b = 6m - 10 b = 12m - 11 now set b = b to get: 6m - 10 = 12m - 11

hero (hero):

Now finish solving for m

OpenStudy (anonymous):

subtract 6m from both sides

hero (hero):

Let me know what m equals when you're done.

OpenStudy (anonymous):

6m+-1

OpenStudy (anonymous):

6m=-1 sorry

hero (hero):

Subtract 6m from both sides; add 11 to both sides to get 11 - 10 = 12m - 6m

OpenStudy (anonymous):

-1/6

hero (hero):

11 - 10 = 1 12m - 6m = 6m

hero (hero):

You end up with 1 = 6m

hero (hero):

then dividing both sides by 6 to get m = ?

OpenStudy (anonymous):

so m= -1/6 right?

hero (hero):

m = 1/6 I was trying help you realize you made a sign mistake.

OpenStudy (anonymous):

Oh I see..don't know why I do that

OpenStudy (anonymous):

so the slope is 1/6

hero (hero):

If you had it right the first time, I would have told you it was correct.

hero (hero):

Anyway, now you know an alternate way to find the slope. Just be careful with the signs.

OpenStudy (anonymous):

Thanks a bunch

OpenStudy (anonymous):

I use to be so good at this haven't done math in forever

OpenStudy (anonymous):

best subject health science and history. Your the best.